Does including a gift with a request for a donation influence the proportion who will make...

  1. Does including a gift with a request for a donation influence the proportion who will make a donation? This question was investigated in a recent study. Letters were sent to a large number of potential donors in Japan. The letter requested a donation for funding schools in Mali. The letter recipients were assigned at random to one of three groups. Those in the first group received a letter with no gift. Those in the second group received a letter that included a small gift (a postcard), and those in the third group received a letter with a larger gift (four postcards). The response of interest was whether or not the letter resulted in a donation.

    Donation No Donation
    No Gift 36 435
    Small Gift 78 502
    Large Gift 72 399


    Perform a test of independence. What is the critical value for the study at a 5% level of significance?

    a.

    5.99

    b.

    1.645

    c.

    11.07

    d.

    7.81

    e.

    3.84
Answer #1
Given table data is as below
MATRIX col1 col2 TOTALS
row 1 36 435 471
row 2 78 502 580
row 3 72 399 471
TOTALS 186 1336 N = 1522
------------------------------------------------------------------

calculation formula for E table matrix
E-TABLE col1 col2
row 1 row1*col1/N row1*col2/N
row 2 row2*col1/N row2*col2/N
row 3 row3*col1/N row3*col2/N
------------------------------------------------------------------

expected frequecies calculated by applying E - table matrix formulae
E-TABLE col1 col2
row 1 57.56 413.44
row 2 70.88 509.12
row 3 57.56 413.44
------------------------------------------------------------------

calculate chisquare test statistic using given observed frequencies, calculated expected frequencies from above
Oi Ei Oi-Ei (Oi-Ei)^2 (Oi-Ei)^2/Ei
36 57.56 -21.56 464.83 8.08
435 413.44 21.56 464.83 1.12
78 70.88 7.12 50.69 0.72
502 509.12 -7.12 50.69 0.1
72 57.56 14.44 208.51 3.62
399 413.44 -14.44 208.51 0.5
ᴪ^2 o = 14.14

------------------------------------------------------------------

set up null vs alternative as
null, Ho: no relation b/w X and Y OR X and Y are independent
alternative, H1: exists a relation b/w X and Y OR X and Y are dependent
level of significance, α = 0.05
from standard normal table, chi square value at right tailed, ᴪ^2 α/2 =5.99
since our test is right tailed,reject Ho when ᴪ^2 o > 5.99
we use test statistic ᴪ^2 o = Σ(Oi-Ei)^2/Ei
from the table , ᴪ^2 o = 14.14
critical value
the value of |ᴪ^2 α| at los 0.05 with d.f (r-1)(c-1)= ( 3 -1 ) * ( 2 - 1 ) = 2 * 1 = 2 is 5.99
we got | ᴪ^2| =14.14 & | ᴪ^2 α | =5.99
make decision
hence value of | ᴪ^2 o | > | ᴪ^2 α| and here we reject Ho
ᴪ^2 p_value =0


ANSWERS
---------------
null, Ho: no relation b/w X and Y OR X and Y are independent
alternative, H1: exists a relation b/w X and Y OR X and Y are dependent
test statistic: 14.14
critical value: 5.99

option:A
p-value:0
decision: reject Ho

we have enough evidence to support the claim that there is association of The letter requested a donation for funding schools in Mali. The letter recipients were assigned at random to one of three groups